OpenStudy (osanseviero):

Is there a limit in this progression? 1.1 + 2.1 + 2 + 3.1 + 2 + 3 + 4

OpenStudy (osanseviero):

Infinite? and how to demostrate it?

hartnn (hartnn):

which progression is the sequence in ?

OpenStudy (osanseviero):

The question: 1. Determine if the next sequence has a limit (demostrate it) and determine it (hint:find the general term) c) 1,1 + 2,1 + 2 + 3,1 + 2 + 3, + 4...

OpenStudy (osanseviero):

I cant find any order...

hartnn (hartnn):

each term is just the sum of '1st n natural numbers'

hartnn (hartnn):

which is n(n+1)/2

hartnn (hartnn):

thats your general term \(a_n =n(n+1)/2\)

hartnn (hartnn):

how do you find whether next sequence has a limit ?

OpenStudy (osanseviero):

Why is that the general term?

OpenStudy (osanseviero):

a1=1(1+1)/2=2/2=1 Does not apply...

hartnn (hartnn):

1st term is indeed 1 right ? 1, 1+2 , 1+2+3 , .... 1,3,6, 10,...

hartnn (hartnn):

and each term is sum of 1st n natural numbers 1st term = 1 2nd term = sum of 1st 2 natural numbers =1+2 =3 and so on

OpenStudy (osanseviero):

Oh...now I understand! Give me a minute

OpenStudy (osanseviero):

Cant you add the series plus iteslf? like this: Sn= 1 + 3 + 6 + 10 + ... + n-2 + n-1 + n Sn= n + n-1 + n-2 +...+ 6 + 3 + 1

OpenStudy (osanseviero):

So: 2Sn= (n+1) + (n+2) + (n+4)...ehm...this is not going right

hartnn (hartnn):

so you want to find the sum ?

hartnn (hartnn):

\(\Large (1/2)\sum (n^2+n) = (1/2)\sum n^2+(1/2)\sum n=...?\)

OpenStudy (osanseviero):

Oh...I thought that I could add Sn + Sn, and then divide them by two to get the general term

hartnn (hartnn):

possible but i don't think its that easy....

OpenStudy (osanseviero):

then how to get the limit?

hartnn (hartnn):

how do you find whether next sequence has a limit ? i may know it, but not in these "limit of next sequence" terms ...

OpenStudy (osanseviero):

When it says "the next sequence" it refers to the one I put

hartnn (hartnn):

ok, so you just want to find whether the sum converges or not ?

OpenStudy (osanseviero):

Look, I have 5 exercises 1. \[\left( -1 \right)^{n}\left( \frac{ 5n+4 }{ 2n } \right)\] For this one we saw that the 4 will be nothing in comparison so the limit without the -1n is 5/2. It has no limit because it is ocilating. 2. \[2^{-1}\times2^{2}\times2^{-3}...\] Because there is a negative sign, it will change, so there is no limit

OpenStudy (osanseviero):

What means converges?

hartnn (hartnn):

sum converge means sum = finite

OpenStudy (osanseviero):

I want to know if it has a limit and to demostrate and determine it. I think that for this infinite is the limit, am I right?

hartnn (hartnn):

yes, it is infinite

OpenStudy (osanseviero):

Thanks :)

OpenStudy (osanseviero):

How did you wrote the general term? Is there a mathematical way to get it?

hartnn (hartnn):

the sum of 1st n natural numbers ? that you can find by your method, Sn = 1+2+3+....n Sn = n+n-1 +n-2 +...3+2+1 2Sn = (1+n)+ (1+n) +..... so, each pair sum is (1+n) and there are n terms so, 2Sn = n (n+1)

hartnn (hartnn):

so, sum of 1st n natural numbers = n (n+1)/2

hartnn (hartnn):

that became your general term and for your sequence, you need to sum this, and whne n becomes large the sum goes on increaseing which menas this progression has no limit.

hartnn (hartnn):

sorry for the typos :P

OpenStudy (osanseviero):

Thanks for a lot

hartnn (hartnn):

welcome ^_^

OpenStudy (osanseviero):

\[an?\left\{ an=\frac{ 1 }{ 2 }\left( a _{n-1} + \frac{ 10 }{ a _{n-1} }\right) \right\} for n \ge2\]

OpenStudy (osanseviero):

that is an =

hartnn (hartnn):

have you solved such problems, with a (n-1) , before.... because i just could think of finding, a1,a2,a3...and see whether there's any pattern...

OpenStudy (osanseviero):

ok, let me try

OpenStudy (osanseviero):

but I dont have a1

OpenStudy (osanseviero):

this is just telling you that the new value is the last one added by 10 divded by iteslf \[\left( a ^{2} +\frac{ 10 }{ a ^{2} }\right)\] divided by two

hartnn (hartnn):

yes, i though it was already given....

OpenStudy (osanseviero):

nope :/

OpenStudy (osanseviero):

So can it be determined with the things given?

hartnn (hartnn):

i am not sure about this....can you ask in new post so that others can try.....sorry.